# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v2

@@ Abstract
Facial analysis is a key component in a wide range of applications such as security, autonomous driving, entertainment, and healthcare. Despite the availability of various facial RGB datasets, the thermal modality, which plays a crucial role in life sciences, medicine, and biometrics, has been largely overlooked. To address this gap, we introduce the T-FAKE dataset, a new large-scale synthetic thermal dataset with sparse and dense landmarks. To facilitate the creation of the dataset, we propose a novel RGB2Thermal loss function, which enables the transfer of thermal style to RGB faces. By utilizing the Wasserstein distance between thermal and RGB patches and the statistical analysis of clinical temperature distributions on faces, we ensure that the generated thermal images closely resemble real samples. Using RGB2Thermal style transfer based on our RGB2Thermal loss function, we create the T-FAKE dataset, a large-scale synthetic thermal dataset of faces. Leveraging our novel T-FAKE dataset, probabilistic landmark prediction, and label adaptation networks, we demonstrate significant improvements in landmark detection methods on thermal images across different landmark conventions. Our models show excellent performance with both sparse 70-point landmarks and dense 478-point landmark annotations. Our code and models are available at https://github.com/phflot/tfake.
*Equal contribution.11footnotetext: Systems Neuroscience & Neurotechnology Unit, Faculty of Medicine, Saarland University & htw saar22footnotetext: Institute of Mathematics, Technische Universität Berlin33footnotetext: Max Planck Institute for Informatics, Saarland Informatics Campus
@@ 1 Introduction
Functional thermal imaging is an acquisition modality with growing importance in many areas from life sciences over medicine to biometrics. Due to the close relation of body temperature with physiological states, thermal imaging of faces emerged as an important tool for screening infectious diseases [54, 46, 4] and for the analysis of neurocognitive and affective states [25]. As a consequence, there exists a growing need for thermal images with annotated facial landmarks. However, in contrast to RGB images, annotated data for thermal images of faces is still rarely available. For example, the largest thermal datasets with facial landmark annotations contain only tens to hundreds of individuals [34, 13, 52, 3], while the largest datasets for face detection for RGB images reached the mark of a million faces (of more than 500k individuals) with the MegaFace benchmark already in 2016 [29]. Moreover, large synthetic RGB datasets, like FAKE [61], appeared to be valuable for sparse and dense landmark detection training.
Generally, the detection of facial landmarks is an important pre-processing task for further facial analysis. These 2D coordinates of corresponding points across faces are part of methods, e.g., for alignment of faces to extract regions of interest [62], the animation of 3D avatars [40], motion magnification [15, 14] or deepfakes [55]. The most common convention for facial landmark annotations uses 68 points across the face and face boundary, see, e.g. [58]. Even though the 68-point convention remains challenging for thermal images [51, 42], there is additional demand for dense landmarkers with more than 300 points. [61, 62, 40].
@@
Faces in thermal images share structural similarities with RGB images. However, temporally varying environmental factors such as temperature and rain droplets, and physiological factors such as blood perfusion changes and vascular details lead to much higher variability in the appearance of faces within individuals [22]. Sweating, flushing, or outdoor rain can result in time-dependent texture changes. Features that are very stable in RGB images such as the contrast of the nose and the eyes can become inverted when the nose or the face is cooled down. Due to the structural similarities, landmarkers trained on RGB datasets were also used under specific thermal imaging conditions [13]. However, they show high failure rates and produce imprecise landmark predictions on challenging thermal images. Up to now, there exist only a few sparse thermal landmark detection methods [34, 42, 52, 38, 37], and we are unaware of dense thermal or a multimodal RGB+Thermal landmarkers.
@@
In this paper, we aim to address this gap by introducing the T-FAKE dataset, the first large-scale synthetic thermal dataset featuring both sparse and dense landmarks for facial analysis. We propose to thermalize a synthetic RGB dataset using our novel RGB2Thermal loss function, which comprises three key components: i) a supervised data term that controls the generation of thermal faces based on a small subset of RGB-thermal pairs, ii) a Wasserstein distance term that aligns the patch distributions of the generated synthetic thermal images with those of real thermal images, and iii) a term incorporating prior information on clinical temperature statistics for different facial regions. This loss enables our model to generalize beyond available lab-condition data to ‘in-the-wild’ images. Figure 4 shows T-FAKE samples and illustrates the advantage of our semi-supervised model over a supervised model trained solely with lab-condition images. To validate the quality of our T-FAKE dataset, we employ a probabilistic landmark prediction method by minimizing the negative log-likelihood function of the landmarks which are assumed to follow a Gaussian distribution as proposed in [62]. Moreover, by introducing a learnable adapter between different landmark conventions, we unify various methods and datasets for evaluation.
@@
We thoroughly evaluate the landmark prediction method trained on our T-FAKE dataset on the established RGB dataset 300W [58] and the challenging thermal dataset CHARLOTTE [3], which includes controlled modulation of temperature, pose, and resolution. Our model demonstrates superior performance on the thermal landmark prediction benchmark compared to previous methods while maintaining performance on the RGB images comparable to state-of-the-art RGB methods. Fig. 2 illustrates the performance of our model for sparse and dense landmarking. In summary, we provide the following contributions:
@@
  • The first large-scale facial synthetic thermal dataset T-FAKE with sparse and dense landmark annotations;
@@
  • Training of the up-to-date first dense thermal landmarker and a state-of-the-art multimodal RGB+Thermal sparse landmarker in combination with highly structured benchmarking across different landmark conventions and modalities.
@@ 2 Related Work
Facial RGB landmarking. Different annotation conventions exist for labeled landmarking datasets. An important distinction is between 2D annotations and 3DA-2D annotations which project 3D landmarks onto images [11]. 2D is closer to human perception, but 3DA-2D enables annotation of occluded face regions. The most widely used annotation convention defines 68-point 2D landmarks available for many facial alignment datasets [19, 32, 67, 6]. Sometimes points on the eye or other details are additionally added [61, 3]. While an in-depth review of facial RGB landmarking is beyond the scope of this work, we introduce those that are relevant to thermal landmarking methods. Kazemi et al. [28] introduced a popular traditional computer vision method for landmark prediction via the dlib library [30]. Kowalski et al. [35] proposed the deep alignment network (DAN) which uses convolutional neural networks and iteratively refines predictions. Around the same time, Bulat et al. [6] proposed the facial alignment network (FA). Both methods predict heat maps of landmark positions Explicit integration of uncertainty quantification into a landmark prediction model via the LUVLi loss has been proposed by Kumar et al. [36]. Recently, image annotation quality became a limiting factor which led to the emergence of synthetic data with high-quality ground truth [44]. Wood et al. [61, 62] trained 3DA-2D landmarkers with excellent performance using their proposed synthetic FAKE dataset with 100k frames of at least 2k individuals.
@@
Table 1: Comparison of thermal landmark datasets. Number of frames (F) and number of subjects (S) are indicated as well as distance to the camera if known. Unavailable data are marked by a question mark.
@@
Thermal landmarking. Many tailored approaches for thermal landmarking use similar architectures as RGB landmarkers and are mainly developed alongside the publication of new datasets, see Table 1. Kuzdeuov et al. [37] released the TFW dataset with with 5 facial landmarks and bounding boxes and trained YOLOv5 networks for landmark prediction. The SF-TL54 dataset [38] has been used for 2D landmark prediction with a dlib and a U-net model. Kopaczka et al. [34] also use a dlib model for their AACHEN dataset with 68-point 2D landmarks. A highly structured 2D landmarking dataset is the CHARLOTTE thermal face dataset with controlled temperature changes by Ashrafi et al [3] It contains around 10k images in total and 7684 images with annotations. In contrast to these publicly available datasets, datasets with limited or no public availability include the the ARL-VTF dataset [52] and a dataset recorded in SARS-CoV-2 drive-in stations (DRIVE-IN) [13] (six minutes recordings at $50$Hz of 436 subjects). The authors of the DRIVE-IN dataset found that a stack of pre-processing functions already leads to a good performance of existing RGB landmark models on thermal images. Similarly, Poster et al. [53] proposed to use transfer learning on a siamese DAN to improve thermal landmark prediction with RGB facial landmarks. Mallat et al. [42] used RGB2Thermal image translation (‘thermalization’), active appearance models, and deep alignment networks of RGB datasets for thermal landmark prediction. In contrast to the previous work, our dataset contains thermal images of at least 2k individuals that are annotated with sparse 70 3DA-2D as well as dense 478 2D point landmarks.
@@
Thermalization. In addition to thermal landmark datasets, various datasets of paired and aligned thermal and RGB images without annotations exist. These can be employed for the training of thermalization models. In contrast to the rich literature on Thermal2RGB synthesis [56, 64, 45], the literature on RGB2Thermal image synthesis using such datasets is more limited. Mallat et al. thermalized facial images by minimizing a perceptual loss [42] using the VIS-TH dataset [41]. Similarly, generative adversarial networks (GANs) have been used for Thermal2RGB synthesis [63, 22] using the NCKU-VTF [22] dataset and the ARL-MMDF [63] dataset containing facial image data. In contrast to the limited public availability of these two datasets, a publicly available dataset with paired facial images is the multi-modal disguise Sejong face database (SEJONG) with 100 subjects, 15 disguise variations and 4849 RGB+Thermal image pairs by Cheema and Moon [7]. Note that all these datasets and resulting models are mostly restricted to frontal views of faces with controlled lighting. Thermalization of general scenes has been achieved by Kniaz et al. [31] with a modified Pix2Pix GAN, where they used the publicly available ThermalWorld dataset with 5098 triplets of RGB images, thermal images, and segmentation masks. In contrast to previous works, we employ a semi-supervised RGB2Thermal loss using a semantic segmentation to improve generalization beyond lab-condition frontal views to ‘in-the-wild’ images.
@@ 3 T-FAKE Dataset
Recently, the success of facial landmarkers trained solely on the FAKE dataset [61], synthetic facial landmarking dataset, has demonstrated the potential of synthetic data for the task. We aim to transfer this success story to the largely overlooked thermal domain and the multimodal setting by introducing our synthetic thermal landmarking dataset T-FAKE featuring at least 2k unique subjects. In particular, we train a semi-supervised image translation model for the thermalization of RGB images using real and synthetic data. To account for data variations in terms of poses, shadows, and expressions, we employ a semi-supervised loss.
@@ 3.1 Thermalization
To generate our large-scale synthetic T-FAKE dataset, we propose a thermalization framework. First, we train our network in a semi-supervised manner based on training pairs of RGB+Thermal images $(\mathbf{X}_{\text{ RGB}},\mathbf{X}_{\text{T}})=(X_{\text{ RGB}}^{n},X_{\text{T}}^{n})_{n=1}^{N}$ from the lab-recorded SEJONG dataset. Further, we use the synthetic RGB images from FAKE $\mathbf{X}_{\text{FAKE}}=(X_{\text{FAKE}}^{m})_{m=1}^{M}$ for which no thermal counterpart exits. The thermalization of these data appears to be difficult due to the large differences between real-world lab-recorded and synthetic in-the-wild data when facing thermalization tasks. For RGB2Thermal translation, we optimize a U-net [57] $T_{\theta}$ by minimizing the loss function
$\mathcal{L}(\theta)$ $:=\frac{1}{N}\sum_{n=1}^{N}{\|T_{\theta}(X_{\text{ RGB}}^{n})-X_{\text{T}}^{n}\|}_{2}^{2}$ (1)
@@
This loss allows us to simultaneously train a domain-adaptive image translation model from paired and unpaired data. The first summand (1) penalizes the deviation of the generated thermal image from the true one and serves as a reconstruction loss for known image pairs. The second summand (2) exploits the Wasserstein distance on the patch distribution of thermalized FAKE images and natural thermal images. This style loss ensures that a generated image displays the thermal image statistics. To be more precise, recall that the squared Wasserstein-2 distance $\mathcal{W}_{2}^{2}$ between empirical measures $\mu=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$ and $\nu=\frac{1}{L}\sum_{l=1}^{L}\delta_{y_{l}}$ for Dirac probability measures $\delta_{\bullet}$ is given by
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\pi\in\Pi}F(\pi):=\sum_{k=1}^{K}\sum_{l=1}^{L}\|x_{k}-y_{l}\|^{2}\pi_{k,l},$ (4)
@@
$\mathcal{W}_{2,E}^{2}(\mu,\nu)$ $:=\inf_{\pi\in\Pi}\big{\{}F(\pi)+\lambda_{E}\sum_{k=1}^{K}\sum_{l=1}^{L}\pi_{k,l}\log\left(\pi_{k,l}\right)\big{\}}$ (5)
with some (small) regularization $\lambda_{E}>0$ has been proposed [9]. This term and its gradients can efficiently be calculated via the Sinkhorn algorithm [12, 50]. To use this loss for thermal translation, we focus on the patch distributions of the generated and real images. For that purpose, we now extract (overlapping) image patches by defining the patch extraction operator $P_{i}X:=X[i_{1}:i_{1}+s,i_{2}:i_{2}+s]$, $i=(i_{1},i_{2})\in I$ of size $s\times s$ and considering the empirical measures generated by the patch distributions of $T_{\theta}(\mathbf{X}_{\text{FAKE}})$,
$\mu_{\text{FAKE}}:=\tfrac{1}{{\color[rgb]{0,0,0}M}}\tfrac{1}{|I|}\sum_{{\color[rgb]{0,0,0}m}=1}^{{\color[rgb]{0,0,0}M}}\sum_{i=1}^{|I|}\delta_{P_{i}\left(T_{\theta}(X_{\text{FAKE}}^{n})\right)},$ (6)
@@
$\operatorname{W}(T_{\theta}(\mathbf{X}_{\text{FAKE}}),\mathbf{X}_{\text{T}})$ $:=\mathcal{W}_{2,E}^{2}(\mu_{\text{FAKE}},\mu_{\text{T}}).$ (7)
During training, we apply a multiscale approach by applying this regularizer over multiple image scales and summing over the results (see Table 4). As a result, the patch statistics of downscaled images are incorporated into the empirical loss. The proposed patch distribution matching-based regularization was inspired by its application in related tasks such as image reconstruction [1, 50], style transfer [39], texture reconstruction [20, 2], and image translation [26].
@@
where the weights $\omega_{i}$ are proportional to the region sizes and $\overline{S_{i}}$ is the mean temperature in the $i$-th region. Note that facial temperatures depend on the environmental conditions. Hence, we train $T_{\theta}$ two times with two different reference temperature sets to simulate ‘cold’ and ‘warm’ environmental conditions. The different regularizers are shown in Fig. 3 (left) and the segmentations in Fig. 4 (bottom).
Predictions of the normalized temperature with our model compared to a model without regularization are shown in Fig. 4. The advantages of our regularization in a domain-adapted setting are clearly visible. The use of regularization reduces image artifacts. The patch regularizer (2) promotes natural texture synthesis and because of our segmentation regularizer (3) temperature differences between facial areas are more pronounced. Hence, our method allows for the adaptation of images with shadows, various skin colors, and distinctive facial expressions. In particular, such features may be underrepresented in facial thermal imaging datasets recorded in lab conditions. Note that the resulting domain adaptation needs for synthetic data solved by a semi-supervised approach based on the access to ground truth segmentations differentiates our task from the RGB2Thermal tasks in [31, 42] and (supervised) Thermal2RGB tasks in [64, 45, 22]. The generalization limitations of supervised models are further discussed in Section 5.5, see Table 4, and the supplementary.

### v3

@@ Abstract
Facial analysis is a key component in a wide range of applications such as healthcare, autonomous driving, and entertainment. Despite the availability of various facial RGB datasets, the thermal modality, which plays a crucial role in life sciences, medicine, and biometrics, has been largely overlooked. To address this gap, we introduce the T-FAKE dataset, a new large-scale synthetic thermal dataset with sparse and dense landmarks. To facilitate the creation of the dataset, we propose a novel RGB2Thermal loss function, which enables the domain-adaptive transfer of RGB faces to thermal style. By utilizing the Wasserstein distance between thermal and RGB patches and the statistical analysis of clinical temperature distributions on faces, we ensure that the generated thermal images closely resemble real samples. Using RGB2Thermal style transfer based on our RGB2Thermal loss function, we create the large-scale synthetic thermal T-FAKE dataset with landmark and segmentation annotations. Leveraging our novel T-FAKE dataset, probabilistic landmark prediction, and label adaptation networks, we demonstrate significant improvements in landmark detection methods on thermal images across different landmark conventions. Our models show excellent performance with both sparse 70-point landmarks and dense 478-point landmark annotations. Moreover, our RGB2Thermal loss leads to notable results in terms of perceptual evaluation and temperature prediction.
@@ 1 Introduction
Functional thermal imaging is an acquisition modality with growing importance in many areas from life sciences over medicine to biometrics. *Equal Contribution; https://github.com/phflot/tfake Due to the close relation of body temperature with physiological states, thermal imaging of faces emerged as an important tool for screening infectious diseases [53, 45, 4] and for the analysis of neurocognitive and affective states [24]. As a consequence, there exists a growing need for thermal images with annotated facial landmarks. However, in contrast to RGB images, annotated data for thermal images of faces is still rarely available. For example, the largest thermal datasets with facial landmark annotations contain only tens to hundreds of individuals [33, 13, 51, 3], while the largest datasets for face detection for RGB images reached the mark of a million faces (of more than 500k individuals) with the MegaFace benchmark already in 2016 [28]. Moreover, large synthetic RGB datasets, like FAKE [60], appeared to be valuable for sparse and dense landmark detection training. Generally, the detection of facial landmarks is an important task for further facial analysis. These 2D coordinates of corresponding points across faces are relevant for, e.g., face alignment [61], animation of 3D avatars [39], motion magnification [15, 14] or deepfakes [54]. The most common convention for facial landmark annotations uses 68 points across the face and face boundary [57]. Even though the 68-point convention remains challenging for thermal images [50, 41], there is additional demand for dense landmarkers with more than 300 points. [60, 61, 39].
@@
Faces in thermal images share structural similarities with RGB images. However, temporally varying environmental factors such as temperature and rain droplets, and physiological factors such as blood perfusion changes and vascular details lead to much higher variability in the appearance of faces within individuals [21]. Sweating, flushing, or outdoor rain can result in time-dependent texture changes. Features that are very stable in RGB images such as the contrast of the nose and the eyes can become inverted when the nose or the face is cooled down. Thus, RGB landmarkers display decent results on thermal images [13], but show high failure rates and imprecise landmark predictions on challenging thermal images. Up to now, there exist only a few sparse thermal landmark detection methods [33, 41, 51, 37, 36], and we are unaware of dense thermal or a multimodal RGB+Thermal landmarkers.
@@
In this paper, we aim to address this gap by introducing the T-FAKE dataset, the first large-scale synthetic thermal dataset featuring both sparse and dense landmarks for facial analysis. We propose to thermalize a synthetic RGB dataset using our novel RGB2Thermal loss function, which comprises three key components: i) a supervised data term that controls the generation of thermal faces based on a small subset of RGB-thermal pairs, ii) a Wasserstein distance term that aligns the patch distributions of the generated synthetic thermal images with those of real thermal images, and iii) a term incorporating prior information on clinical temperature statistics for different facial regions. This loss enables our model to generalize beyond available lab-condition data to ‘in-the-wild’ images. Figure 2 shows T-FAKE samples and illustrates the advantage of our semi-supervised model over a supervised model trained solely with lab-condition images. To validate the quality of our T-FAKE dataset, we employ a probabilistic landmark prediction method with integrated face detection by minimizing the negative log-likelihood function of the landmarks which are assumed to follow a Gaussian distribution as proposed in [61]. Moreover, by introducing a learnable adapter between different landmark conventions, we unify various methods and datasets for evaluation.
@@
We thoroughly evaluate the landmark prediction method trained on our T-FAKE dataset on the challenging thermal dataset CHARLOTTE [3], which includes controlled modulation of temperature, pose, and resolution and on the established RGB dataset 300W [57]. Our model demonstrates superior performance on the thermal landmark prediction benchmark compared to previous methods while maintaining performance on RGB images comparable to state-of-the-art RGB methods. Figure 1 illustrates the performance of our model for sparse and dense landmarking.
In summary, we provide the following contributions:
@@
  • The first large-scale facial synthetic thermal dataset T-FAKE with 70 sparse and 478 dense landmarks and additional semantic segmentation masks;
@@
  • Training of the up-to-date first dense thermal landmarker and a state-of-the-art multimodal RGB+Thermal sparse landmarker in combination with highly structured benchmarking across landmark conventions and modalities.
@@ 2 Related Work
Facial RGB landmarking. Different annotation conventions exist for labeled landmarking datasets. An important distinction is between 2D annotations and 3DA-2D annotations which project 3D landmarks onto images [11]. 2D is closer to human perception, but 3DA-2D enables annotation of occluded face regions. The most widely used annotation convention defines 68-point 2D landmarks available for many facial alignment datasets [18, 31, 66, 6]. Sometimes points on the eye or other details are additionally added [60, 3]. While an in-depth review of facial RGB landmarking is beyond the scope of this work, we introduce those that are relevant to thermal landmarking methods. Kazemi et al. [27] introduced a popular traditional computer vision method for landmark prediction via the dlib library [29]. Kowalski et al. [34] proposed the deep alignment network (DAN) which uses convolutional neural networks and iteratively refines predictions. Around the same time, Bulat et al. [6] proposed the facial alignment network (FA). Both methods predict heat maps of landmark positions. Explicit integration of uncertainty quantification into a landmark prediction model via the LUVLi loss has been proposed by Kumar et al. [35]. Recently, image annotation quality became a limiting factor which led to the emergence of synthetic data with high-quality ground truth [43]. Wood et al. [60, 61] trained 3DA-2D landmarkers with excellent performance using their proposed synthetic FAKE dataset with 100k frames of at least 2k individuals. Note that most landmarkers are trained on facial bounding boxes. As a result, they are either evaluated in combination with a face detector or on pre-defined ground truth bounding boxes. Here, we focus on the combined evaluation of face and landmark detection, a harder, but more practical task.
@@
Table 1: Thermal landmark dataset comparison. Indicated are the number of frames (F) and subjects (S), camera distance, and if recorded outside of a controlled laboratory. Unavailable data are marked by ✗ and ✗ indicates no response received from corresponding authors.
@@
Thermal landmarking. Many tailored approaches for thermal landmarking use similar architectures as RGB landmarkers and are mainly developed alongside the publication of new datasets, see Table 1. Kuzdeuov et al. [36] released the TFW dataset with with 5 facial landmarks and bounding boxes and trained YOLOv5 networks for landmark prediction. The SF-TL54 dataset [37] has been used for 2D landmark prediction with a dlib and a U-net model. Kopaczka et al. [33] use a DAN for their AACHEN dataset with 68-point 2D landmarks. A highly structured 2D landmarking dataset is the CHARLOTTE thermal face dataset with controlled temperature changes by Ashrafi et al [3]. It contains around 10k images in total and 7684 images with annotations. In contrast to these publicly available datasets, datasets with limited or no public availability include the the ARL-VTF dataset [51] and a dataset recorded in SARS-CoV-2 drive-in stations (DRIVE-IN) [13] (six minutes recordings at $50$Hz of 436 subjects). The authors of the DRIVE-IN dataset found that a stack of pre-processing functions already leads to a good performance of existing RGB landmark models on thermal images. Note that such thermal datasets are often captured with uncooled microbolometers with automatic radiometric calibration. This leads to comparable temperature values but may introduce small measurement errors. Beyond thermal images, Poster et al. [52] proposed to use transfer learning on a siamese DAN to improve thermal landmark prediction with RGB facial landmarks. Mallat et al. [41] used RGB2Thermal image translation (‘thermalization’), active appearance models, and deep alignment networks of RGB datasets for thermal landmark prediction. In contrast to the previous work, our dataset contains thermal images of at least 2k individuals that are annotated with sparse 70 3DA-2D and dense 478 2D point landmarks.
@@
Thermalization. In addition to thermal landmark datasets, various datasets of paired and aligned thermal and RGB images without annotations exist. These can be employed for the training of thermalization models. In contrast to the rich literature on Thermal2RGB synthesis [55, 63, 44], the literature on RGB2Thermal image synthesis using such datasets is more limited. Mallat et al. thermalized facial images by minimizing a perceptual loss [41] using the VIS-TH dataset [40]. Similarly, generative adversarial networks (GANs) have been used for Thermal2RGB synthesis [62, 21] using the NCKU-VTF [21] dataset and the ARL-MMDF [62] dataset containing facial image data. In contrast to the limited public availability of these two datasets, a publicly available dataset with paired facial images is the multi-modal disguise Sejong face database (SEJONG) with 100 subjects, 15 disguise variations and 4849 RGB+Thermal image pairs by Cheema and Moon [7]. Note that all these datasets and resulting models are mostly restricted to frontal views of faces with controlled lighting. Thermalization of general scenes has been achieved by Kniaz et al. [30] with a modified Pix2Pix GAN based on the 5098 triplets of RGB images, thermal images, and segmentation masks in the ThermalWorld dataset. In contrast to previous works, we employ a semi-supervised RGB2Thermal loss using a semantic segmentation to improve generalization beyond lab-condition frontal views to ‘in-the-wild’ images.
@@ 3 T-FAKE Dataset
Recently, the success of facial landmarkers trained solely on the FAKE dataset [60], a synthetic facial landmarking dataset, has demonstrated the potential of synthetic data for the task. We aim to transfer this success story to the largely overlooked thermal domain and the multimodal setting by introducing our synthetic thermal landmarking dataset T-FAKE featuring at least 2k unique subjects. In particular, we train a semi-supervised image translation model for the thermalization of RGB images using real and synthetic data. To account for data variations in terms of poses, shadows, and expressions, we employ a semi-supervised loss.
@@ 3.1 Thermalization
To generate our large-scale synthetic T-FAKE dataset, we propose a thermalization framework (see Figure 3). First, we train our network in a semi-supervised manner based on training pairs of RGB+Thermal images $(\mathbf{X}_{\text{ RGB}},\mathbf{X}_{\text{T}})=(X_{\text{ RGB}}^{n},X_{\text{T}}^{n})_{n=1}^{N}$ from the lab-recorded SEJONG dataset. Further, we use the synthetic RGB images from FAKE $\mathbf{X}_{\text{FAKE}}=(X_{\text{FAKE}}^{m})_{m=1}^{M}$ for which no thermal counterpart exits. The thermalization of these data appears to be difficult due to the large differences between real-world lab-recorded and synthetic in-the-wild data when facing thermalization tasks. For RGB2Thermal translation, we optimize a U-net [56] $T_{\theta}$ by minimizing the loss function
$\mathcal{L}(\theta)$ $\coloneqq\frac{1}{N}\sum_{n=1}^{N}{\|T_{\theta}(X_{\text{ RGB}}^{n})-X_{\text{T}}^{n}\|}_{2}^{2}$ (1)
@@
This loss allows us to train a supervised image translation model from paired data while ensuring domain-generalization of the model using unpaired data. The first summand (1) penalizes the deviation of the generated thermal image from the true one and serves as a reconstruction loss for known image pairs. The second summand (2) exploits the Wasserstein distance on the patch distribution of thermalized FAKE images and natural thermal images. This style loss ensures that a generated image displays the thermal image statistics. To be more precise, recall that the squared Wasserstein-2 distance $\mathcal{W}_{2}^{2}$ between empirical measures $\mu=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$ and $\nu=\frac{1}{L}\sum_{l=1}^{L}\delta_{y_{l}}$ for Dirac probability measures $\delta_{\bullet}$ is given by
$\mathcal{W}_{2}^{2}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi}F(\pi)\coloneqq\inf_{\pi\in\Pi}\sum_{k=1}^{K}\sum_{l=1}^{L}\|x_{k}-y_{l}\|^{2}\pi_{k,l},$ (4)
@@
$\mathcal{W}_{2,E}^{2}(\mu,\nu)$ $\coloneqq\inf_{\pi\in\Pi}\big{\{}F(\pi)+\lambda_{E}\sum_{k=1}^{K}\sum_{l=1}^{L}\pi_{k,l}\log\left(\pi_{k,l}\right)\big{\}}$ (5)
with some (small) regularization $\lambda_{E}>0$ has been proposed [9]. This term and its gradients can efficiently be calculated via the Sinkhorn algorithm [12, 49]. To use this loss for thermal translation, we focus on the (subsampled) patch distributions of the generated and real images. For that purpose, we now extract (overlapping) image patches by defining the patch extractor $P_{i}~{}X~{}\coloneqq~{}X[i_{1}:~{}i_{1}+s,~{}i_{2}~{}:~{}i_{2}~{}+~{}s]$, $i=(i_{1},i_{2})\in I$ of size $s\times s$ and considering the empirical measures generated by the patch distributions of $T_{\theta}(\mathbf{X}_{\text{FAKE}})$,
$\mu_{\text{FAKE}}\coloneqq\tfrac{1}{{M}}\tfrac{1}{|I|}\sum_{{m}=1}^{{M}}\sum_{i=1}^{|I|}\delta_{P_{i}\left(T_{\theta}(X_{\text{FAKE}}^{n})\right)},$ (6)
@@
$\operatorname{W}(T_{\theta}(\mathbf{X}_{\text{FAKE}}),\mathbf{X}_{\text{T}})$ $\coloneqq\mathcal{W}_{2,E}^{2}(\mu_{\text{FAKE}},\mu_{\text{T}}).$ (7)
During training, we apply a multiscale approach by applying this regularizer over multiple image scales and summing over the results. As a result, the patch statistics of downscaled images are incorporated into the empirical loss. The proposed patch distribution matching-based regularization was inspired by its application in related tasks such as image reconstruction [1, 49], style transfer [38], texture reconstruction [19, 2], and image translation [25].
@@
where the weights $\omega_{i}$ are proportional to the region sizes and $\overline{S_{i}}$ is the mean temperature in the $i$-th region. Note that facial temperatures depend on the environmental conditions. Hence, we train $T_{\theta}$ two times with two different reference temperature sets to simulate ‘cold’ and ‘warm’ environmental conditions. The segmentations are shown in Figure 2 (bottom) and different regularizers in Figure 3 (left).
Predictions of the normalized temperature with our model compared to a model without regularization are shown in Figure 2. The advantages of our regularization in a domain-adapted setting are clearly visible. The use of regularization reduces image artifacts. The patch regularizer (2) promotes natural texture synthesis and because of our segmentation regularizer (3) temperature differences between facial areas are more pronounced. Hence, our method allows for the adaptation of images with shadows, various skin colors, and distinctive facial expressions. In particular, such features may be underrepresented in facial thermal imaging datasets recorded in lab conditions. Note that the resulting domain adaptation needs for synthetic data solved by a semi-supervised approach based on the access to ground truth segmentations differentiates our task from the RGB2Thermal tasks in [30, 41] and (supervised) Thermal2RGB tasks in [63, 44, 21]. Generalization limitations of supervised models are discussed in the supplementary.